OpenStudy (anonymous):

Solve the initial value problem using Laplace transforms...\[y''+5y'-6y=21e^{t-1}\space ;\space y(1)=-1 \space , \space y'(1)=9\]

OpenStudy (anonymous):

\[y''+5y'-6y=21e^{t-1}\space ;\space y(1)=-1 \space , \space y'(1)=9\]

OpenStudy (anonymous):

I have a feeling I need to do a time shift

OpenStudy (anonymous):

Here's another probably easier one with the same type of setup...\[y''-y=t-2 \space ; \space y(2)=3 \space , \space y'(2)=0\]

OpenStudy (anonymous):

That one might be easier to go thru on here